# Non-linear degenerate parabolic flow equations and a finer differential structure on Wasserstein spaces

### v1

@@ 3.6.2 First Main Result: Generalised Central Limit Theorem—weak version
$\qquad\biggl{|}\int_{M}\varphi\operatorname{d}(\mu^{N}(\mu^{N}(x_{l^{N}(s)}))-\mu^{N}(0))$- (3.93)
$\sum_{i=0}^{l^{N}(s)-1}\biggl{[}\int_{x_{i}}^{x_{i+1}}\int_{M}\int_{\mathcal{X}(M)}\tfrac{1}{2}\mathcal{L}^{2}_{\overline{X}[\mu^{N}(x_{i})]}\varphi(x)\operatorname{d}V[\mu^{N}(x_{i})](X)\operatorname{d}\mu^{N}(x_{i})(x)\operatorname{d}t$
$+\int_{x_{i}}^{x_{i+1}}\int_{M}\mathcal{L}_{\overline{V}[\mu^{N}\bigl{(}\frac{x_{i}+x_{i+1}}{2}\bigr{)}]}(\varphi)\operatorname{d}\mu^{N}\bigl{(}\frac{x_{i}+x_{i+1}}{2}\bigr{)}(x)\operatorname{d}t\biggr{]}\biggr{|}$
$\qquad\leq C\|\varphi\|_{\mathcal{C}^{3}}x_{i_{l^{N}(s)}}(\max_{i}\bigl{(}x_{i+1}-x_{i}\bigr{)})^{\frac{\min(1,p-2)}{2}}\leq C\|\varphi\|_{\mathcal{C}^{3}}x_{i_{l^{N}(s)}}\frac{1}{N^{\frac{\min(1,p-2)}{2}}}.$
@@
$\sum_{i=0}^{l^{N}(s)-1}\biggl{[}\int_{x_{i}}^{x_{i+1}}\int_{M}\int_{\mathcal{X}(M)}\tfrac{1}{2}\mathcal{L}^{2}_{\overline{X}[\mu^{N}(x_{i})]}\varphi(x)\operatorname{d}V[\mu^{N}(x_{i})](X)\operatorname{d}\mu^{N}(x_{i})(x)\operatorname{d}t$ (3.94)
$+\int_{x_{i}}^{x_{i+1}}\int_{M}\mathcal{L}_{\overline{V}[\mu^{N}\bigl{(}\frac{x_{i}+x_{i+1}}{2}\bigr{)}]}(\varphi)\operatorname{d}\mu^{N}\bigl{(}\frac{x_{i}+x_{i+1}}{2}\bigr{)}(x)\operatorname{d}t\biggr{]}$
$=\int_{0}^{x_{i_{l^{N}(s)}}}\sum_{i=0}^{l^{N}(s)-1}\int_{M}\int_{\mathcal{X}(M)}{\mathbf{1}}_{[x_{i},x_{i+1}]}(t)\bigl{(}\tfrac{1}{2}\mathcal{L}^{2}_{\overline{X}[\mu^{N}(x_{i})]}\varphi(x)\bigr{)}$
$\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\operatorname{d}V[\mu^{N}(x_{i})](X)\operatorname{d}\mu^{N}(x_{i})(x)\operatorname{d}t$
$+\int_{0}^{x_{i_{l^{N}(s)}}}\sum_{i=0}^{l^{N}(s)-1}\int_{M}{\mathbf{1}}_{[x_{i},x_{i+1}]}(t)\bigl{(}\mathcal{L}_{\overline{V}[\mu^{N}\bigl{(}\frac{x_{i}+x_{i+1}}{2}\bigr{)}]}(\varphi)\bigr{)}\operatorname{d}\mu^{N}\bigl{(}\frac{x_{i}+x_{i+1}}{2}\bigr{)}(x)\operatorname{d}t.$
@@
$\sum_{i=0}^{l^{N}(s)-1}$ $\biggl{[}\int_{x_{i}}^{x_{i+1}}\int_{M}\int_{\mathcal{X}(M)}\tfrac{1}{2}\mathcal{L}^{2}_{\overline{X}[\mu^{N}(t)]}\varphi(x)\operatorname{d}V[\mu^{N}(t)](X)\operatorname{d}\mu^{N}(t)(x)\operatorname{d}t$ (3.95)
@@
$=\int_{0}^{x_{i_{l^{N}(s)}}}\int_{M}\int_{\mathcal{X}(M)}\tfrac{1}{2}\mathcal{L}^{2}_{\overline{X}[\mu^{N}(t)]}\varphi(x)\operatorname{d}V[\mu^{N}(t)](X)\operatorname{d}\mu^{N}(t)(x)\operatorname{d}t$
$+\int_{0}^{x_{i_{l^{N}(s)}}}\int_{M}\tfrac{1}{2}\mathcal{L}_{\overline{V}[\mu^{N}(t)]}\operatorname{d}\mu^{N}(t)(x)\operatorname{d}t.$
@@
Let $\varepsilon>0$. Since the pp–th moments of $V$ are uniformly bounded, they are tight in the 22–nd moment by Lemma 2.11. Hence, there exists $R>0$ such that for all $\mu\in\mathcal{W}_{p}(M)$,
$\int_{\mathcal{X}^{2,\infty}\setminus B(0,R)}\|X\|_{W^{2,\infty}}^{2}\,\operatorname{d}V[\mu(t)](X)<\varepsilon.$ (3.96)
@@
$\biggl{|}$ $\int_{\mathcal{X}(M)\setminus B(0,R)}\tfrac{1}{2}\mathcal{L}^{2}_{\overline{X}[\mu^{N}(t)]}\varphi(x)-\psi(X)\operatorname{d}(V[\mu^{N}(t)]-V[\mu^{N}(x_{i})])(X)\biggr{|}$ (3.97)
$\,\leq\int_{\mathcal{X}(M)\setminus B(0,R)}\bigl{|}\tfrac{1}{2}\mathcal{L}^{2}_{\overline{X}[\mu^{N}(t)]}\varphi(x)-\psi(X)\bigr{|}\operatorname{d}(V[\mu^{N}(t)]+V[\mu^{N}(x_{i})])(X)$
$\,\leq\int_{\mathcal{X}(M)\setminus B(0,R)}\tfrac{1}{2}\|X\|_{W^{2,\infty}}^{2}(\|\nabla\varphi\|_{\infty}+\|\nabla^{2}\varphi\|_{\infty})\operatorname{d}(V[\mu^{N}(t)]+V[\mu^{N}(x_{i})])(X)$
@@
On the other hand, for all $x\in M$, by the Monge–Kantorovich-duality 2.3,
$\biggl{|}\int_{\mathcal{X}(M)}$ $\psi(X)\operatorname{d}(V[\mu^{N}(t)]-V[\mu^{N}(x_{i})])(X)\biggr{|}$ (3.98)
$\leq CR(\|\nabla\varphi\|_{\infty}+\|\nabla^{2}\varphi\|_{\infty})\mathcal{W}_{1}\bigl{(}V[\mu^{N}(t)],V[\mu^{N}(x_{i})]\bigr{)}$
Now, $V$ is uniformly continuous, and for all $\mathbb{N}\in\mathbb{N}$, $\mathcal{W}_{p}(\mu^{N}(t),\mu^{N}{x_{i_{l}^{N}(t)}})\leq\mathcal{B}\sqrt{x_{i_{l}^{N}(t)+1}-x_{i_{l}^{N}(t)}}$. Hence,
@@
$\biggl{|}\int_{\mathcal{X}(M)}\psi(X)\operatorname{d}(V[\mu^{N}(t)]$ $-V[\mu^{N}(x_{i_{l}^{N}(t)})])(X)\biggr{|}$ (3.99)
@@
Thus, the function $x\mapsto\int_{\mathcal{X}(M)}\tfrac{1}{2}\mathcal{L}^{2}_{\overline{X}[\mu^{N}(t)]}\varphi(x)\operatorname{d}(V[\mu^{N}(t)]-V[\mu^{N}(x_{i_{l}^{N}(t)})])(X)$ uniformly converges to 0 for $N\to\infty$. This implies that there exists a constant $K_{1}>0$ such that
$\biggl{|}\int_{0}^{x_{i_{l^{N}(s)}}}\sum_{i=0}^{I_{N}-1}\int_{M}\int_{\mathcal{X}(M)}{\mathbf{1}}_{[x_{i},x_{i+1}]}(t)\frac{1}{2}\mathcal{L}^{2}_{\overline{X}[\mu^{N}(t)]}\varphi(x)$ (3.100)
$\qquad\qquad\operatorname{d}(V[\mu^{N}(x_{i})](X)-V[\mu^{N}(t)])(X)\operatorname{d}\mu^{N}(x_{i})(x)\operatorname{d}t\biggr{|}$
$\qquad\qquad\qquad\qquad\qquad\qquad\qquad\leq K_{1}(\|\nabla\varphi\|_{\infty}+\|\nabla^{2}\varphi\|_{\infty})|s|\omega^{V}(\frac{1}{\sqrt{N}}),$
@@
$\biggl{|}$ $\int_{\mathcal{X}(M)}\tfrac{1}{2}\bigl{(}\mathcal{L}^{2}_{\overline{X}[\mu^{N}(t)]}\varphi(x)-\mathcal{L}^{2}_{\overline{X}[\mu^{N}(t)]}\varphi(y)\bigr{)}\operatorname{d}V[\mu^{N}(t)](X)\biggr{|}$ (3.101)
$\quad\leq C^{\prime}\int_{\mathcal{X}(M)}d(x,y)\|X\|_{W^{2,\infty}}^{2}(\|\nabla\varphi\|_{\infty}+\|\nabla^{2}\varphi\|_{\infty}+\|\nabla^{3}\varphi\|_{\infty})\operatorname{d}V[\mu^{N}(t)](X)$
$\quad\leq C^{\prime}(\|\nabla\varphi\|_{\infty}+\|\nabla^{2}\varphi\|_{\infty}+\|\nabla^{3}\varphi\|_{\infty})(B+1)d(x,y).$
Hence, $x\mapsto\int_{\mathcal{X}(M)}\frac{1}{2}\mathcal{L}^{2}_{\overline{X}[\mu^{N}(t)]}\varphi(x)\operatorname{d}V[\mu^{N}(t)](X)$ is Lipschitz with constant $K_{2}(\|\nabla\varphi\|_{\infty}+\|\nabla^{2}\varphi\|_{\infty}+\|\nabla^{3}\varphi\|_{\infty})$ for some $K_{2}>0$. Again, Lemma 2.11, the Monge–Kantorovic duality and the proof of Proposition 2.4 yield
$\biggl{|}\int_{0}^{x_{i_{l^{N}(s)}}}$ $\sum_{i=0}^{l-1}\int_{M}\int_{\mathcal{X}(M)}{\mathbf{1}}_{[x_{i},x_{i+1}]}(t)\frac{1}{2}\mathcal{L}^{2}_{\overline{X}[\mu^{N}(t)]}\varphi(x)$ (3.102)
@@
$\qquad\leq K_{2}\mathcal{B}(\|\nabla\varphi\|_{\infty}+\|\nabla^{2}\varphi\|_{\infty}+\|\nabla^{3}\varphi\|_{\infty})\frac{|s|}{\sqrt{N}}.$
@@
Thus, the last point of Lemma 3.29 and a similar argument as the those above yields that there exists $K_{3}>0$ such that
$\biggl{|}\int_{0}^{x_{i_{l^{N}(s)}}}$ $\sum_{i=0}^{l-1}\int_{M}\int_{\mathcal{X}(M)}{\mathbf{1}}_{[x_{i},x_{i+1}]}(t)\frac{1}{2}(\mathcal{L}^{2}_{\overline{X}[\mu^{N}(t)]}-\mathcal{L}^{2}_{\overline{X}[\mu^{N}(x_{i})]})\varphi(x)$ (3.104)
@@
$\quad\leq K_{3}(\|\nabla\varphi\|_{\infty}+\|\nabla^{2}\varphi\|_{\infty})\widetilde{\omega}^{V}(\frac{\mathcal{B}}{\sqrt{N}})|s|.$
@@
$\biggl{|}$ $\int_{0}^{s}\int_{M}\int_{\mathcal{X}(M)}\frac{1}{2}\mathcal{L}^{2}_{\overline{X}[\mu^{N}(t)]}\varphi(x)\operatorname{d}V[\mu^{N}(t)](X)\operatorname{d}\mu^{N}(t)(x)\operatorname{d}t$ (3.105)
$\qquad-\sum_{i=0}^{l^{N}(s)-1}\int_{x_{i}}^{x_{i+1}}\int_{M}\int_{\mathcal{X}(M)}\frac{1}{2}\mathcal{L}^{2}_{\overline{X}[\mu^{N}(x_{i})]}\varphi(x)$
$\qquad\qquad\qquad\qquad\qquad\qquad\operatorname{d}V[\mu^{N}(x_{i})](X)\operatorname{d}\mu^{N}(x_{i})(x)\operatorname{d}t\biggr{|}$
$\qquad\qquad\qquad\qquad\leq K_{4}(\|\nabla\varphi\|_{\infty}+\|\nabla^{2}\varphi\|_{\infty}+\|\nabla^{3}\varphi\|_{\infty})\omega(\frac{\mathcal{B}}{\sqrt{N}})|s|.$
@@
$-\int_{0}^{s}\int_{M}\mathcal{L}_{\overline{X}_{V[\mu(t)]}}(\varphi)\operatorname{d}\mu(t)\operatorname{d}t\biggr{|}$
$\qquad\qquad\qquad\qquad\qquad\qquad\qquad\leq K_{5}(\|\nabla\varphi\|_{\infty}+\|\nabla^{2}\varphi\|_{\infty})\frac{\mathcal{B}}{\sqrt{N}}|s|.$
@@
$-\int_{0}^{s}\int_{M}\int_{\mathcal{X}(M)}\frac{1}{2}\mathcal{L}^{2}_{\overline{X}[\mu^{N}(t)]}\varphi(x)\operatorname{d}V[\mu^{N}(t)](X)\operatorname{d}\mu^{N}(t)(x)\operatorname{d}t$
@@
$\qquad\leq\mathcal{K}(\|\nabla\varphi\|_{\infty}+\|\nabla^{2}\varphi\|_{\infty}+\|\nabla^{3}\varphi\|_{\infty})\Omega(\frac{1}{N})|s|.$
for some oncreasing continuous $\Omega$ with $\Omega(0)=0$, which concludes the second step.

### v2

@@ 3.6.5 First Main Result: Generalised Central Limit Theorem—weak version
$\biggl|\int_{M}\varphi\operatorname{d}(\mu^{N}(\mu^{N}(x_{l^{N}(s)}))-\mu^{N}(0))$ (3.99)
$-\sum_{i=0}^{l^{N}(s)-1}\biggl[\int_{x_{i}}^{x_{i+1}}\int_{M}\int_{\mathcal{X}^{2,\infty}(M)}\tfrac{1}{2}\mathcal{L}^{2}_{\overline{X}[\mu^{N}(x_{i})]}\varphi(x)\operatorname{d}V[\mu^{N}(x_{i})](X)\operatorname{d}\mu^{N}(x_{i})(x)\operatorname{d}t$
$+\int_{x_{i}}^{x_{i+1}}\int_{M}\mathcal{L}_{\overline{V}\bigl[\mu^{N}\bigl(\frac{x_{i}+x_{i+1}}{2}\bigr)\bigr]}(\varphi)\operatorname{d}\mu^{N}\bigl(\frac{x_{i}+x_{i+1}}{2}\bigr)(x)\operatorname{d}t\biggr]\biggr|$
$\hskip 18.49988pt\leq C\|\varphi\|_{\mathcal{C}^{3}}x_{i_{l^{N}(s)}}(\max_{i}\bigl(x_{i+1}-x_{i}\bigr))^{\frac{\min(1,p-2)}{2}}\leq C\|\varphi\|_{\mathcal{C}^{3}}x_{i_{l^{N}(s)}}\frac{1}{N^{\frac{\min(1,p-2)}{2}}}.$
@@
$\sum_{i=0}^{l^{N}(s)-1}\biggl[\int_{x_{i}}^{x_{i+1}}\int_{M}\int_{\mathcal{X}^{2,\infty}(M)}\tfrac{1}{2}\mathcal{L}^{2}_{\overline{X}[\mu^{N}(x_{i})]}\varphi(x)\operatorname{d}V[\mu^{N}(x_{i})](X)\operatorname{d}\mu^{N}(x_{i})(x)\operatorname{d}t$ (3.100)
$\hskip 18.49988pt+\int_{x_{i}}^{x_{i+1}}\int_{M}\mathcal{L}_{\overline{V}\bigl[\mu^{N}\bigl(\frac{x_{i}+x_{i+1}}{2}\bigr)\bigr]}(\varphi)\operatorname{d}\mu^{N}\bigl(\frac{x_{i}+x_{i+1}}{2}\bigr)(x)\operatorname{d}t\biggr]$
$=\int_{0}^{x_{i_{l^{N}(s)}}}\sum_{i=0}^{l^{N}(s)-1}\int_{M}\int_{\mathcal{X}^{2,\infty}(M)}{\mathbf{1}}_{[x_{i},x_{i+1}]}(t)\bigl(\tfrac{1}{2}\mathcal{L}^{2}_{\overline{X}[\mu^{N}(x_{i})]}\varphi(x)\bigr)$
$\hskip 18.49988pt\hskip 18.49988pt\hskip 18.49988pt\hskip 18.49988pt\hskip 18.49988pt\hskip 18.49988pt\hskip 18.49988pt\hskip 18.49988pt\hskip 18.49988pt\operatorname{d}V[\mu^{N}(x_{i})](X)\operatorname{d}\mu^{N}(x_{i})(x)\operatorname{d}t$
$+\int_{0}^{x_{i_{l^{N}(s)}}}\sum_{i=0}^{l^{N}(s)-1}\!\!\int_{M}{\mathbf{1}}_{[x_{i},x_{i+1}]}(t)\biggl(\mathcal{L}_{\overline{V}\bigl[\mu^{N}\bigl(\frac{x_{i}+x_{i+1}}{2}\bigr)\bigr]}(\varphi)\biggr)\operatorname{d}\mu^{N}\bigl(\frac{x_{i}+x_{i+1}}{2}\bigr)(x)\operatorname{d}t.$
@@
$\sum_{i=0}^{l^{N}(s)-1}$ $\biggl[\int_{x_{i}}^{x_{i+1}}\int_{M}\int_{\mathcal{X}^{2,\infty}(M)}\tfrac{1}{2}\mathcal{L}^{2}_{\overline{X}[\mu^{N}(t)]}\varphi(x)\operatorname{d}V[\mu^{N}(t)](X)\operatorname{d}\mu^{N}(t)(x)\operatorname{d}t$ (3.101)
@@
$=\int_{0}^{x_{i_{l^{N}(s)}}}\int_{M}\int_{\mathcal{X}^{2,\infty}(M)}\tfrac{1}{2}\mathcal{L}^{2}_{\overline{X}[\mu^{N}(t)]}\varphi(x)\operatorname{d}V[\mu^{N}(t)](X)\operatorname{d}\mu^{N}(t)(x)\operatorname{d}t$
$+\int_{0}^{x_{i_{l^{N}(s)}}}\int_{M}\mathcal{L}_{\overline{V}[\mu^{N}(t)]}\operatorname{d}\mu^{N}(t)(x)\operatorname{d}t.$
@@
Let $\varepsilon>0$. Since the pp–th moments of $V$ are uniformly bounded, they are tight in the 22nd moment by Lemma 2.18. Hence, there exists $R>0$ such that for all $\mu\in\mathcal{W}_{p}(M)$ and for $\mathcal{X}^{2,\infty}_{\geq R}(M):=\mathcal{X}^{2,\infty}(M)\setminus B(0,R)$,
$\int_{\mathcal{X}^{2,\infty}_{\geq R}(M)}\|X\|_{W^{2,\infty}}^{2}\,\operatorname{d}V[\mu(t)](X)<\varepsilon.$ (3.102)
@@
$\biggl|$ $\int_{\mathcal{X}^{2,\infty}_{\geq R}(M)}\tfrac{1}{2}\mathcal{L}^{2}_{\overline{X}[\mu^{N}(t)]}\varphi(x)-\psi(X)\operatorname{d}(V[\mu^{N}(t)]-V[\mu^{N}(x_{i})])(X)\biggr|$ (3.103)
$\,\leq\int_{\mathcal{X}^{2,\infty}_{\geq R}(M)}\bigl|\tfrac{1}{2}\mathcal{L}^{2}_{\overline{X}[\mu^{N}(t)]}\varphi(x)-\psi(X)\bigr|\operatorname{d}(V[\mu^{N}(t)]+V[\mu^{N}(x_{i})])(X)$
$\,\leq\int_{\mathcal{X}^{2,\infty}_{\geq R}(M)}\tfrac{1}{2}\|X\|_{W^{2,\infty}}^{2}(\|\nabla\varphi\|_{\infty}+\|\nabla^{2}\varphi\|_{\infty})\operatorname{d}(V[\mu^{N}(t)]+V[\mu^{N}(x_{i})])(X)$
@@
On the other hand, for all $x\in M$, by the Monge–Kantorovich-duality 2.10,
$\biggl|\int_{\mathcal{X}^{2,\infty}(M)}$ $\psi(X)\operatorname{d}(V[\mu^{N}(t)]-V[\mu^{N}(x_{i})])(X)\biggr|$ (3.104)
$\leq CR(\|\nabla\varphi\|_{\infty}+\|\nabla^{2}\varphi\|_{\infty})\mathcal{W}_{1}\bigl(V[\mu^{N}(t)],V[\mu^{N}(x_{i})]\bigr).$
Now, $V$ is uniformly continuous, and for all $N\in\mathbb{N}$, $\mathcal{W}_{p}(\mu^{N}(t),\mu^{N}{x_{i_{l}^{N}(t)}})\leq\mathcal{B}\sqrt{x_{i_{l}^{N}(t)+1}-x_{i_{l}^{N}(t)}}$. Hence,
@@
$\biggl|\int_{\mathcal{X}^{2,\infty}(M)}\psi(X)\operatorname{d}(V[\mu^{N}(t)]$ $-V[\mu^{N}(x_{i_{l}^{N}(t)})])(X)\biggr|$ (3.105)
@@
Thus, the function
$x\mapsto\int_{\mathcal{X}^{2,\infty}(M)}\tfrac{1}{2}\mathcal{L}^{2}_{\overline{X}[\mu^{N}(t)]}\varphi(x)\operatorname{d}(V[\mu^{N}(t)]-V[\mu^{N}(x_{i_{l}^{N}(t)})])(X)$
uniformly converges to 0 for $N\to\infty$. This implies that there exists a constant $K_{1}>0$ such that
$\biggl|\int_{0}^{x_{i_{l^{N}(s)}}}\sum_{i=0}^{I_{N}-1}\int_{M}\int_{\mathcal{X}^{2,\infty}(M)}{\mathbf{1}}_{[x_{i},x_{i+1}]}(t)\frac{1}{2}\mathcal{L}^{2}_{\overline{X}[\mu^{N}(t)]}\varphi(x)$ (3.106)
$\hskip 18.49988pt\hskip 18.49988pt\operatorname{d}(V[\mu^{N}(x_{i})](X)-V[\mu^{N}(t)])(X)\operatorname{d}\mu^{N}(x_{i})(x)\operatorname{d}t\biggr|$
$\hskip 18.49988pt\hskip 18.49988pt\hskip 18.49988pt\hskip 18.49988pt\hskip 18.49988pt\hskip 18.49988pt\hskip 18.49988pt\leq K_{1}(\|\nabla\varphi\|_{\infty}+\|\nabla^{2}\varphi\|_{\infty})|s|\omega^{V}(\frac{1}{\sqrt{N}}),$
@@
$\biggl|$ $\int_{\mathcal{X}^{2,\infty}(M)}\tfrac{1}{2}\bigl(\mathcal{L}^{2}_{\overline{X}[\mu^{N}(t)]}\varphi(x)-\mathcal{L}^{2}_{\overline{X}[\mu^{N}(t)]}\varphi(y)\bigr)\operatorname{d}V[\mu^{N}(t)](X)\biggr|$ (3.107)
$\qquad\leq C^{\prime}\int_{\mathcal{X}^{2,\infty}(M)}d(x,y)\|X\|_{W^{2,\infty}}^{2}(\|\nabla\varphi\|_{\infty}+\|\nabla^{2}\varphi\|_{\infty}+\|\nabla^{3}\varphi\|_{\infty})$
$\hskip 18.49988pt\hskip 18.49988pt\hskip 18.49988pt\hskip 18.49988pt\hskip 18.49988pt\hskip 18.49988pt\hskip 18.49988pt\hskip 18.49988pt\hskip 18.49988pt\cdot\operatorname{d}V[\mu^{N}(t)](X)$
$\qquad\leq C^{\prime}(\|\nabla\varphi\|_{\infty}+\|\nabla^{2}\varphi\|_{\infty}+\|\nabla^{3}\varphi\|_{\infty})(B+1)d(x,y).$
Hence,
$x\mapsto\int_{\mathcal{X}^{2,\infty}(M)}\frac{1}{2}\mathcal{L}^{2}_{\overline{X}[\mu^{N}(t)]}\varphi(x)\operatorname{d}V[\mu^{N}(t)](X)$
is Lipschitz with constant $K_{2}(\|\nabla\varphi\|_{\infty}+\|\nabla^{2}\varphi\|_{\infty}+\|\nabla^{3}\varphi\|_{\infty})$ for some $K_{2}>0$. Again, Lemma 2.18, the Monge–Kantorovic duality and the proof of Proposition 2.11 yield
$\biggl|\int_{0}^{x_{i_{l^{N}(s)}}}$ $\sum_{i=0}^{l-1}\int_{M}\int_{\mathcal{X}^{2,\infty}(M)}{\mathbf{1}}_{[x_{i},x_{i+1}]}(t)\frac{1}{2}\mathcal{L}^{2}_{\overline{X}[\mu^{N}(t)]}\varphi(x)$ (3.108)
@@
$\hskip 18.49988pt\leq K_{2}\mathcal{B}(\|\nabla\varphi\|_{\infty}+\|\nabla^{2}\varphi\|_{\infty}+\|\nabla^{3}\varphi\|_{\infty})\frac{|s|}{\sqrt{N}}.$
@@
Thus, the last point of Lemma 3.31 and a similar argument as those above yield that there exists $K_{3}>0$ such that
$\biggl|\int_{0}^{x_{i_{l^{N}(s)}}}$ $\sum_{i=0}^{l-1}\int_{M}\int_{\mathcal{X}^{2,\infty}(M)}{\mathbf{1}}_{[x_{i},x_{i+1}]}(t)\frac{1}{2}(\mathcal{L}^{2}_{\overline{X}[\mu^{N}(t)]}-\mathcal{L}^{2}_{\overline{X}[\mu^{N}(x_{i})]})\varphi(x)$ (3.110)
@@
$\qquad\leq K_{3}(\|\nabla\varphi\|_{\infty}+\|\nabla^{2}\varphi\|_{\infty})\widetilde{\omega}^{V}(\frac{\mathcal{B}}{\sqrt{N}})|s|.$
@@
$\biggl|$ $\int_{0}^{s}\int_{M}\int_{\mathcal{X}^{2,\infty}(M)}\frac{1}{2}\mathcal{L}^{2}_{\overline{X}[\mu^{N}(t)]}\varphi(x)\operatorname{d}V[\mu^{N}(t)](X)\operatorname{d}\mu^{N}(t)(x)\operatorname{d}t$ (3.111)
$\hskip 18.49988pt-\sum_{i=0}^{l^{N}(s)-1}\int_{x_{i}}^{x_{i+1}}\int_{M}\int_{\mathcal{X}^{2,\infty}(M)}\frac{1}{2}\mathcal{L}^{2}_{\overline{X}[\mu^{N}(x_{i})]}\varphi(x)$
$\hskip 18.49988pt\hskip 18.49988pt\hskip 18.49988pt\hskip 18.49988pt\hskip 18.49988pt\hskip 18.49988pt\operatorname{d}V[\mu^{N}(x_{i})](X)\operatorname{d}\mu^{N}(x_{i})(x)\operatorname{d}t\biggr|$
$\hskip 18.49988pt\hskip 18.49988pt\hskip 18.49988pt\hskip 18.49988pt\leq K_{4}(\|\nabla\varphi\|_{\infty}+\|\nabla^{2}\varphi\|_{\infty}+\|\nabla^{3}\varphi\|_{\infty})\omega(\frac{\mathcal{B}}{\sqrt{N}})|s|.$
@@
$\hskip 18.49988pt\hskip 18.49988pt-\int_{0}^{s}\int_{M}\mathcal{L}_{\overline{X}_{V[\mu(t)]}}(\varphi)\operatorname{d}\mu(t)\operatorname{d}t\biggr|$
$\hskip 18.49988pt\hskip 18.49988pt\hskip 18.49988pt\hskip 18.49988pt\hskip 18.49988pt\hskip 18.49988pt\hskip 18.49988pt\leq K_{5}(\|\nabla\varphi\|_{\infty}+\|\nabla^{2}\varphi\|_{\infty})\frac{\mathcal{B}}{\sqrt{N}}|s|.$
@@
$-\int_{0}^{s}\int_{M}\int_{\mathcal{X}^{2,\infty}(M)}\frac{1}{2}\mathcal{L}^{2}_{\overline{X}[\mu^{N}(t)]}\varphi(x)\operatorname{d}V[\mu^{N}(t)](X)\operatorname{d}\mu^{N}(t)(x)\operatorname{d}t$
@@
$\hskip 18.49988pt\leq\mathcal{K}(\|\nabla\varphi\|_{\infty}+\|\nabla^{2}\varphi\|_{\infty}+\|\nabla^{3}\varphi\|_{\infty})\Omega(\tfrac{1}{N})|s|.$
for some nondecreasing continuous $\Omega$ with $\Omega(0)=0$, which concludes the second step.